OpenStudy (anonymous):

Coordinates for D. (?,6) C is (10,0) Rectangle ABCD. A(2,14) B(-2,8) C(10,0) D(?,6)

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

earlier we found line BC to be \[\Large y = -\frac{2}{3}x+\frac{20}{3}\]

OpenStudy (anonymous):

Yes, would it be the same for AD?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

but the same steps will be used find the slope of AD (since it's parallel to BC, the two slopes are equal). Then use it to find the equation of line AD

jimthompson5910 (jim_thompson5910):

you also need the equation for line CD

OpenStudy (anonymous):

For AD equation, would it start like: y-14=m(x-2)? m = slope

jimthompson5910 (jim_thompson5910):

yes and m = -2/3 because AD and BC have the same slope (they are parallel lines)

OpenStudy (anonymous):

So, then I can plug in 6 to get the x value for D?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

you'll need equation CD

jimthompson5910 (jim_thompson5910):

oh wait, nvm

jimthompson5910 (jim_thompson5910):

yes you can plug in y = 6. I didn't see D(?,6)

OpenStudy (anonymous):

Yeah, sorry, with my packet, I have image of it showing the C point being on 0. So, helped me determine the distance for the AD y and what to subtract.

OpenStudy (anonymous):

-2/3 * -2 would be 12/3?

jimthompson5910 (jim_thompson5910):

-2/3 * -2 = 4/3

OpenStudy (anonymous):

6-14=-2/3(x-2) -8=-2/3x + 4/3 It ends up being: 6 = -2/3x + 28/3 right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now solve for x

OpenStudy (anonymous):

x=5?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

oh I see what went wrong

jimthompson5910 (jim_thompson5910):

When you went from -8=-2/3x + 4/3 to 6 = -2/3x + 28/3 you made an error the 6 in ` 6 = -2/3x + 28/3` should be 0 so you should have 0 = -2/3x + 28/3

jimthompson5910 (jim_thompson5910):

it's 0 because you added 8 to both sides the -8 and +8 combine to 0 on the left side

OpenStudy (anonymous):

OH. x=14

jimthompson5910 (jim_thompson5910):

yep D is (14,6)

OpenStudy (anonymous):

To be honest, I read my graph wrong aswell, so I was confused. I lack in graphing, but all is well. You really helped and taught me a few things, I hope to remember. (:

jimthompson5910 (jim_thompson5910):

I'm glad I was of help